OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x – y = –4 4x – 3y = –6 Answer 4x – 3(–12x – 4) = –6 4(–12x – 4) – 3y = –6 4(12x + 4) – 3y = –6 4x – 3(12x + 4) = –6

OpenStudy (anonymous):

4(12x + 4) – 3y = –6

OpenStudy (anonymous):

12x - y = -4 ==> add y to both sides and add 4 to both sides=> 12x + 4 = y Substitute in 2nd equation ==> 4x - 3(12x + 4) = -6 Correct answer is the last one